OpenStudy (ellerai):

5. Determine if the lines represented by 2x + 3y = 15 and y = 3/2x – 6 are parallel, perpendicular, neither, or the same line.

OpenStudy (mathstudent55):

What do you know about the slopes of parallel lines?

OpenStudy (ellerai):

that they are positive?

OpenStudy (mathstudent55):

No. Slopes of lines in general can be positive, negative, zero, or even undefined (when there is division by 0). Parallel lines have the same slope. The slopes of parallel lines are equal. Ok?

OpenStudy (ellerai):

yes

OpenStudy (mathstudent55):

Next. What about the slopes of perpendicular lines?

OpenStudy (ellerai):

they are the opposite of the slope

OpenStudy (mathstudent55):

The slopes of perpendicular lines are negative reciprocals. That means if two lines are perpendicular, their slopes have a product of -1. That also means, if you see the slope of a line, then the slope of a perpendicular to that line is obtained by flipping the fraction and changing the sign of the slope of the first line.

OpenStudy (ellerai):

okay

OpenStudy (mathstudent55):

Examples: parallel lines: two lines have the slope 2/3 perpendicular lines: one line has slope 2/3, and the other line has slope -3/2

OpenStudy (ellerai):

yeah

OpenStudy (mathstudent55):

Now let's look at your problem. Look at the equation of the second line \(y = \dfrac{3}{2} x - 6\) This equation is very convenient for our problem because it is in the slope-intercept form. When you write the equation of a line in the slope-intercept form it looks like this: \(y = mx + b\), and m is the slope. You can read directly from your second equation that the slope of the second line is 3/2.

OpenStudy (ellerai):

yep :)

OpenStudy (mathstudent55):

We already know what the slope of the second line is. Now we need to find out what the slope f the first line is to see if it is the same as this one, making the lines parallel, or if it is the negative reciprocal of this one, making the lines perpendicular, or if it is neither.

OpenStudy (mathstudent55):

Let's take the equation of the first line, and put it in the slope-intercept form, \(y =- mx + b\)

OpenStudy (ellerai):

Okay how do we turn it into slope intercept form tho

OpenStudy (mathstudent55):

We need to solve the equation of the first line for y.

OpenStudy (ellerai):

2x + 3y = 15 so would we isolate y?

OpenStudy (ellerai):

by subtracting 2x from both sides?

OpenStudy (mathstudent55):

\(2x + 3y = 15\) We want y alone on the left side. 2x is being added to the left side, so we get rid of the addition of 2x by subtracting 2x from both sides.

OpenStudy (mathstudent55):

Exactly. You're faster than me. Pretty soon, I'll be asking you for math help.

OpenStudy (ellerai):

Lol :P nah im bad at math

OpenStudy (mathstudent55):

\(2x + 3y = 15\) \(2x -2x + 3y = -2x + 15\) \(3y = -2x + 15\) So far, so good. The next step is that we have 3y, but we want just y on the left side. Since the 3 is multiplying y, we divide both sides by 3.

OpenStudy (mathstudent55):

\(3y = -2x + 15\) \(\dfrac{3y}{3} = -\dfrac{2}{3}x + \dfrac{15}{3}\) Now we simplify the fractions.

OpenStudy (ellerai):

y = -2/3x + 5?

OpenStudy (mathstudent55):

\(y = -\dfrac{2}{3}x + 5\) Correct.

OpenStudy (mathstudent55):

Now we have the slope of the second line. What is it?

OpenStudy (ellerai):

-2/3

OpenStudy (mathstudent55):

Correct. Now compare slopes: \(\dfrac{3}{2} \) and \(-\dfrac{2}{3} \). They are clearly not the same. The lines are definitely not parallel. Are they negative reciprocals of each other making the lines perpendicular?

OpenStudy (ellerai):

yes, because 3/2 as a negative reciprocal is -2/3

OpenStudy (mathstudent55):

Correct. When you flip one of the slopes and change its sign you get the other. More formally in math, when you multiply them both together, you get -1. Both of the above make the slopes negative reciprocals, and the lines are perpendicular.

OpenStudy (mathstudent55):

Great job!

OpenStudy (ellerai):

yay! thank you so much! !

OpenStudy (mathstudent55):

You're very welcome.